OpenStudy (anonymous):

In 1980, median family income was about 17000 and in 2000 it was about 43000 a) find the slope of the line passing through the points (1980,17000) and (2000,43000) b)Interpret the slope as a rate of change. c) if this trend continues, estimate the median family income in 2007.

OpenStudy (anonymous):

subtract the first y value from the second y value. Divide it by the first x value subtracted from the second x value.\[slope= (y _{2} - y _{1} ) / (x _{2}- x _{1})\] This is the rate of change ( in units of dollars per year)

OpenStudy (anonymous):

1980-2000/1700-43000

OpenStudy (anonymous):

i got 52100 for 2007 but i think i am second guessing

OpenStudy (anonymous):

points are (as you are told) \((1980,17000), (2000,43000)\) slope is \[\frac{43000-17000}{2000-1980}\] which is a rather silly way to look at it, but it works

OpenStudy (anonymous):

the sane way to look at it is that in 20 years, from 1980 to 2000 income increased by \(26000\)

OpenStudy (anonymous):

per year, that is an increase of \(26000\div 20=1300\) per year

OpenStudy (anonymous):

that answers this question b)Interpret the slope as a rate of change. it means income grew by $1300 per year

OpenStudy (anonymous):

ok but how would you find the rate for 2007

OpenStudy (anonymous):

what is the income in 2000?

OpenStudy (anonymous):

43000

OpenStudy (anonymous):

and if it goes up by 1300 per year, what would it be in 2007?

OpenStudy (anonymous):

52100

OpenStudy (anonymous):

let me check because i just asked, didn't do it

OpenStudy (anonymous):

48200

OpenStudy (anonymous):

you were right the first time

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

\[43000+7\times 1300=52100\] btw you can tell the moron who wrote this question that growth of this kind is exponential, not linear

OpenStudy (anonymous):

yw